# OMoE: Diversifying Mixture of Low-Rank Adaptation by Orthogonal Finetuning

### v1

@@ Appendix F Additional visualization of representations using different methods
Appendix G Additional visualization of representations using different number of experts in OMoE
As illustrated in Fig. 9, Fig. 10 and Fig. 11, the representations under different numbers of experts in OMoE are visualized using a uniform t-SNE transformation for 2D representation reduction. The representations are obtained from layer 8, 24, 32 of LLaMA-2 7B.
@@ Appendix G Additional visualization of representations using different number of experts in OMoE
Figure 9: The representations under different numbers of experts in OMoE are visualized using a uniform t-SNE transformation for 2D representation reduction. The representations are obtained from layer 8 of LLaMA-2 7B.
@@
Figure 10: The representations under different numbers of experts in OMoE are visualized using a uniform t-SNE transformation for 2D representation reduction. The representations are obtained from layer 24 of LLaMA-2 7B.
@@
Figure 11: The representations under different numbers of experts in OMoE are visualized using a uniform t-SNE transformation for 2D representation reduction. The representations are obtained from layer 32 of LLaMA-2 7B.

### v2

@@ Appendix D Additional Visualization of Representations using Different Methods
Appendix E Additional Visualization of Representations using Different Number of Experts in OMoE
As illustrated in Fig. 9, Fig. 10 and Fig. 11, the representations under different numbers of experts in OMoE are visualized using a uniform t-SNE transformation for 2D representation reduction. The representations are obtained from layers 8, 24, and 32 of LLaMA-2 7B.
@@ Appendix E Additional Visualization of Representations using Different Number of Experts in OMoE
Figure 9: The representations under different numbers of experts in OMoE using top-2 routing are visualized via t-SNE. The representations are obtained from layer 8 of LLaMA-2 7B.
@@
Figure 10: The representations under different numbers of experts in OMoE using top-2 routing are visualized via t-SNE. The representations are obtained from layer 24 of LLaMA-2 7B.
@@
Figure 11: The representations under different numbers of experts in OMoE using top-2 routing are visualized via t-SNE. The representations are obtained from layer 32 of LLaMA-2 7B.